An altitude of an equilateral triangle is 6 inches. Find the area of the triangle.


1
Expert's answer
2021-02-03T03:47:09-0500

If an altitude of an equilateral triangle is 6 inches, then the side of this triangle is "a = 2h/\sqrt{3} = 12/\sqrt{3} = 4\sqrt{3}."

The area of the triangle is:

"S = 0.5a×h = 0.5×4\sqrt{3}×6 = 12\sqrt{3}"inches.